OpenStudy (anonymous):

graph a function increasing at (-infinity,-2), decreasing at (2,infinity), and constant at (-2,2)

zepdrix (zepdrix):

Ahhhh! it's cookie mostah! :D man that looks delicious +_+

OpenStudy (anonymous):

yeah. I know

OpenStudy (anonymous):

Increasing over \((-\infty,-2)\) means \(f'(x)>0\) for \(x<-2\). Similarly, decreasing over \((2,\infty)\) means \(f'(x)<0\) for \(x>2\). Constant over \((-2,2)\) means \(f'(x)=0\) for all \(x\in(-2,2)\). Judging by this information, \(f(x)\) can be easily modeled by a piecewise function. Try to see if you can fill in the blanks: \[f(x)=\begin{cases}g(x)&\text{for }x<2\\h(x)&\text{for }-2<x<2\\k(x)&\text{for }x>2\end{cases}\] It doesn't matter what \(f(x)\) is assigned to for \(x=\pm2\) in this case, since it's not required that \(f\) be continuous at these points; only that its components are differentiable. Clearly \(h(x)\) is some constant function, so you can choose any number here. Find some \(g,k\) that work.

OpenStudy (anonymous):

OMG!!! thank you so much. It makes sense now!!!

OpenStudy (anonymous):

You're welcome!